OpenStudy (anonymous):

what is the integral of csc^5 6x cot 6x dx???

OpenStudy (turingtest):

strip out a csc\[\int\csc^5(6x)\cot(6x)dx=\int\csc^4(6x)\cdot\csc(6x)\cot(6x)dx\]now it's an easy u-sub\[u=\csc(6x)\to du=-6\csc(6x)\cot(6x)\]\[-\frac16\int u^4du\]

OpenStudy (anonymous):

thanks for ur help!!

OpenStudy (turingtest):

welcome :D

OpenStudy (anonymous):

if u dont mind, can u help me with one more??

OpenStudy (turingtest):

sure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what is the integral of (x dx) / sin^2 4x^2

OpenStudy (turingtest):

\[\int\frac{xdx}{\sin^2(4x^2)}\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

\[\int\frac{xdx}{\sin^2(4x^2)}\]\[u=4x^2\]do you have any idea what to do next?

OpenStudy (turingtest):

in u substitutions we need u and du... we can't integrate without a differential

OpenStudy (anonymous):

let me see, then du= 8x dx

OpenStudy (turingtest):

right

OpenStudy (anonymous):

1/8 du = x dx

OpenStudy (turingtest):

wait I'm sorry that won't work!

OpenStudy (turingtest):

that would be the right answer though...

OpenStudy (anonymous):

so now what do we do next?

OpenStudy (turingtest):

integration by parts

OpenStudy (anonymous):

what do u mean by that?

OpenStudy (turingtest):

\[\int udv=uv-\int vdu\]

OpenStudy (anonymous):

how do u come up with that?

OpenStudy (turingtest):

it's an integration technique you learn in calc II what level calculus are you in?

OpenStudy (anonymous):

calc I

OpenStudy (anonymous):

are there any way we can move sin from denominator to numerator??

OpenStudy (turingtest):

sort of :D\[\frac{x}{\sin^2(4x^2)}=x\csc^2(4x^2)\]now use the same substitution good idea ;)

OpenStudy (anonymous):

oh really??? it was just my random thought

OpenStudy (turingtest):

think you can do it now?

OpenStudy (anonymous):

let me try, im not sure though, just give me a minute

OpenStudy (anonymous):

oh wait a minute, what do i do with cot

OpenStudy (turingtest):

what do you mean ?

OpenStudy (anonymous):

when i did the u substitution, i got du=-8x csc 4x^2 cot 4x^2 dx

OpenStudy (turingtest):

that's not the sub I showed you...

OpenStudy (turingtest):

u=4x^2 du=8xdx 1/8csc^2udu

OpenStudy (anonymous):

but then we can't take antiderivative with csc ^2

OpenStudy (turingtest):

\[u=4x^2\]\[du=8xdx\]\[\frac18\int\csc^2udu\]oh yes we can, think back somethings derivative is -csc^2

OpenStudy (anonymous):

oh yeah so the antiderivaative would be cot, right??

OpenStudy (turingtest):

yeah, but we're missing the negative sign, so we have to add it\[\frac18\int\csc^2udu=-\frac18\int-\csc^2 udu=-\frac18\cot u\]

OpenStudy (turingtest):

...+C

OpenStudy (anonymous):

so the final answer is -1/8 cot 4x^2

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

oh yes, i always forget the +C

OpenStudy (anonymous):

thanks so much!! that was such a big help

OpenStudy (turingtest):

anytime !

OpenStudy (anonymous):

i think i did this wrong

OpenStudy (anonymous):

when i did the u sub du= -8x csc 4x^2 cot 4x^2, so i substitute all those into the integral x csc^2 4x^2, so the x and csc^2 cancelled out but then we have u cot 4 x^2 left, and we still can't integrate it like that

OpenStudy (turingtest):

how did you get that du ? u=4x^2 du=8xdx you don't include the csc in that substitution

OpenStudy (anonymous):

oh never mind, my bad i made this harder than it should be

OpenStudy (turingtest):

\[\int\frac{xdx}{\sin^2(4x^2)}=\int x\csc^2(4x^2)dx\]\[u=4x^2\to du=8xdx\]\[\frac18\int\csc^2udu=-\frac18\cot u+C=-\frac18\cot(4x^2)+C\]ta dah!

OpenStudy (anonymous):

oh yeah i just got it now, sorry!!

OpenStudy (turingtest):

no prob :D

OpenStudy (anonymous):

im sorry to bother u this much, but i have one more problem, if y+x = sin y cos x, what is dy/dx

OpenStudy (turingtest):

implicit differentiation!

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

taking the derivative of both sides with respect to x we need to remember that we need to use the chain rule on functions of y\[\frac d{dx}f(y)=f'(y)y'\]we also need the product rule on the right part\[y'+1=\cos x\cdot\cos y\cdot y'+\sin y\cdot(-\sin x)\]solve for y'

OpenStudy (anonymous):

but dont we need to separate all y to one side and all x to other side?

OpenStudy (turingtest):

you need to get y' by itself, however you want to go about it

OpenStudy (anonymous):

is that the same thing?

OpenStudy (turingtest):

yeah, I guess so you said something about getting x on one side and y on the other though, so it's not quite that just solve for the variable y', it's algebra

OpenStudy (anonymous):

sorry i got it mixed up with integration

OpenStudy (anonymous):

this one ask for derivative, not antiderivative, well i guess cuz i been doing the the antideriv

OpenStudy (turingtest):

I gave you the result of taking the derivative, you just have to solve for the variable

OpenStudy (anonymous):

oh ok thanks :D

OpenStudy (anonymous):

for final answer i got, (-1-sin x sin y) / (1-cos x cos y)

OpenStudy (anonymous):

thanks for helping me today!!

OpenStudy (turingtest):

you're right! you're also welcome :)